OpenStudy (anonymous):

Solve the simultaneous equations y = −3x + 2 y = 2x2 − 7x − 4 I think the way its set out is confusing me?

OpenStudy (anonymous):

the second one is y=2x^

OpenStudy (lgbasallote):

y = -3x + 2 substitute into the second equation -3x +2 = 2x^2 - 7x - 4 2x^2 + 3x - 7x - 2 - 4 = 0 2x^2 - 4x - 6 = 0 x^2 - 2x - 3 = 0 (x-3)(x+1) = 0 x = 3 x = -1

OpenStudy (lgbasallote):

i think you can solve for y on your own?

OpenStudy (anonymous):

so in the first it ends with x=3, y=-1?

OpenStudy (anonymous):

ah no worries I’ve got it you’ve given the x for both now to solve the y cheers